Find the relative extrema of 𝑓 π‘₯ = 3π‘₯^5 βˆ’ 5π‘₯^2.


1
Expert's answer
2021-12-06T16:47:51-0500

"\text{Relative extrema can only occur at critical points}\\
\text{Given}, \, f(x) = 3x^5 - 5x^2. \, \text{Now,}\\
f'(x) = 15x^4 - 10x \\
\text{At critical points,} \, f'(x) = 0 \\
\text{Hence}, 15x^4 - 10x = 0\\
5x(3x^3 - 2) = 0 \\
\implies x = 0, \, \sqrt[3]{\frac{2}{3}}\\
\text{Put these values into} \, f(x) = 3x^5 - 5x^2. \, \text{Now},\\
f(0) = 0 \, \, \text{and}\, \, f(\sqrt[3]{\frac{2}{3}}) = -2.29.\\
\text{Hence,} \, f(x) \, \text{has a local maximum at} \, x = 0 \\
\text{and local minimum at} \, x = \sqrt[3]{\frac{2}{3}}."